OpenStudy (anonymous):

A line passes through the point negative four comma one and has a slope of two-thirds. If you use this slope to generate another point on the line, what are the coordinates of that point? negative one comma three negative two comma four negative twelve comma two negative eight comma three

OpenStudy (hoblos):

the slope between two points (x1,y1) and (x2,y2) is \[m = \frac{ y2-y1 }{ x2 - x1 }\] here you have the point (-4,1) and slope 2/3 suppose the point we want to find is (x,y) so you have \[\frac{ 2 }{ 3 } = \frac{ y - 1 }{ x+4 }\]

OpenStudy (hoblos):

now try replacing the coordinates of the point of each choice and check them

OpenStudy (hoblos):

are you able to solve it now ?